25*25

25*25
 

Grade:12th pass

12 Answers

Nikhil Upadhyay
8 years ago
625
Anoopam Mishra
8 years ago
625
Athiyaman
8 years ago
You can actually find out the squares of any number ending with 5 easilyby using this method.
For eg you want to find the square of 25 , take the number before 5  in this case 2 and multiply it with 3, we get 6.And we also know the square of  5 is 25.So the answer will be 625.Basically what you do is if you have a number xy5, you take all the numbers before 5 and you multiply it with a number (xy+1).So you should get xy*(xy+1).and then you add 25 to the end of the number.We can find the square of any number ending with 5 using this method.85*85 = (8*9) 25 = 7225.
raju
8 years ago
625
raju
8 years ago
5*5*5*5*=625
charan
8 years ago
25*25=625
RAJU
8 years ago
252
SHANMUKESHWAR
8 years ago
625
KARTHIK
8 years ago
625
OUTPASS
8 years ago
625
pa1
8 years ago
625
SEELA YUVARAJU
8 years ago
625